# BMG-Q: Localized Bipartite Match Graph Attention Q-Learning for Ride-Pooling Order Dispatch

### v1

@@ IV-C GATDDQN for Localized Bipartite Match Graph
Remark 2.
Note that not both spatial and temporal interdependence among agents are encoded within our GATDDQN. As the decision-making process of each agent is modeled as an MDP, we have the Markov Property [12], which asserts that the current state of each agent encapsulates all relevant historical information about this agent, as well as historical interactions between the agent and the environment that are pertinent to future decisions. Therefore, when the agent evaluate other agents’ states with GAT, both spatial and temporal correlations have been considered.
@@ V-A Dynamic ILP via Posterior Score Function
In existing literature, the ILP for the bipartite matching process typically relies on $Q(s,a,\theta)$, such as (11) in Section II. However, this approach encounters two issues: (1) it lacks an exploration mechanism in the bipartite matching process, which restricts the exploration of the vehicle agents’ state space; (2) matching based solely on $Q(s,a,\theta)$ is prone to bias due to variability in the estimates, which can affect the decision-making process. These will potentially lead to suboptimal solutions.
To deal with the above two issues, we propose posterior score function to better integrate the value function of GATDDQN with ILP. The formulation of our posterior score function $S(s_{n,t},g_{n,t},a_{n,t})$ is given in Equation (26) below, with its visualization shown in Figure 4. To effectively balance exploration and exploitation, we implement an $\epsilon$-greedy strategy and introduce a term $S_{\text{explore}}$ during the exploration phase. This term represents the upper bound of the Q-value, $S(s_{n,t},g_{n,t},a_{n,t})$, which is set to a significantly high value (e.g., 100,000) to encourage exploration in exploration stage:
@@
Figure 4: Dynamic ILP with posterior score function.
@@ V-B Training GATDDQN for Large-Scale System
$g_{\text{clip}}$ =$=\begin{cases}g\times\frac{\text{threshold}}{||g||_{2}},&\text{if}||g||_{2}>\text{threshold},\\ g,&\text{otherwise},\end{cases}$
@@
Similar to the policy improvement theorem proved in [57, 58, 59], the gradient clipping enforces the agents to update the policy within a region so as to not over-fit to recent experience tuples and guarantee to improve its policy. Through our further training and validation, we find that the two tricks not only help to stabilize the training process but also make the framework more robust to task variability and parameter change.
@@ V-C Summary of the proposed BMG-Q Framework
The training details of our BMG-Q framework could be found at Algorithm 2. Specifically, after initializing the simulator and GATDDQN in steps 1 through 4, we enter the training phase. To achieve a balance between exploitation and exploration, we perform exploration decay to exploration rate of the bipartite matching process in step 5. This gradual reduction in exploration rate is designed to transition the focus from exploration to exploitation as the learning advances. In steps 9 to 12, similar to the DDQN backbone of ILPDDQN, our GATDDDQN backbone adopts double networks, experience replay, and soft update. Thanks to the localized bipartite match graph topology, graph sampling, and gradient clipping introduced in steps 9 to 11, GATDDQN backbone manages to learn to capture the localized interdependence in very large-scale system with thousands of agents, thus leading to more optimal assignment decisions of the overall BMG-Q framework.
@@
1:  Simulator Initialization: OSRM Router Model[32], Number of Vehicles $N$, Matching Distance $R_{match}$
@@
4:     Initialize: Episode Order Requirements, and Number of Vehicles $N$
@@
11:           Sample $N$ experience tuples $(s,g,a,r,s^{\prime},g^{\prime})$ in $M$ as mini-batch $D$ and use Equation (28) and (29) to update $\theta$.
@@
Figure 5: Overview of the proposed BMG-Q framework for ride-pooling vehicles dispatch. For every decision round, initially users submit orders through a mobile application, and the system updates and sorts these orders alongside vehicle information. Taking into account long-term uncertainties including the intricate interdependence of agents, the GATDDQN network evaluates and dynamically assigns orders to suitable vehicles using ILP. After assignments, vehicle agents execute their actions, with experiences collected for subsequent learning phases of the GATDDQN network. Concurrently, the Open Street Routing Machine updates routes and estimated times of arrival, which are communicated back to the central platform. (Part of icons and map are from [60, 61]).
@@ VI-A Simulation Setup
The simulation environment for this study is based on the public dataset of taxi trips in Manhattan, New York City [15, 31]. This dataset includes detailed information for each trip, such as pickup and dropoff times, origin and destination geo-coordinates, trip distance, and duration. Focusing on peak hours—specifically from 8:00 AM to 10:00 AM—we tailored the training dataset to include data from trips that occurred between 8:00 AM and 8:30 AM on May 4, 2016. During this half-hour period, the average order density reached approximately 275 trips per minute in central Manhattan, totaling about 8,250 orders. For the validation dataset, we similarly extracted data from trips within the same half-hour window but on various days throughout May 2016. We divided Manhattan into 57 zones. This zoning was informed by the distribution of orders and a resolution of 800m x 800m was used, for which a visualization is shown in Figure 6. To serve these demand with a minimum service rate of 85% across all RL methodologies tested in this study, we set the number of ride-pooling agents as 1000 and number of vacant seats of each vehicle as 3. To provide real-time route guidance and estimating the passengers’ onboard time, we employed the OSRM model [32] through docker as our router. The coefficients of reward function is set as ${\beta_{0}}=100$, ${\beta_{1}}=40$, ${\beta_{2}}=5$, ${\beta_{3}}=2$, ${\beta_{5}}=20$, ${thre}=15$, with the aim to encourage agents to pick up more orders but not result in large average detours of passengers. For bipartite match process, we set matching distance $R_{match}$ as 1.2 km and any requests that remain unmatched with vehicles for more than five minutes will be automatically rejected. Additionally, it’s important to note that our focus is on fully optimizing the potential of the ride-pooling fleet order dispatch during peak hours, when the demand is high and the occupancy of the vehicles are naturally high. Therefore, we have chosen not to include rebalancing operations [3] in our approaches or any of the benchmark approaches considered in this study. The incorporation of rebalancing operations is left for future work.
@@
In implementing GATDDQN, we adopted a linear transformation as message passing layer and multi-head attentions with a head count of 3 as aggregation layer to form the backbone structure of GATs. For the sake of a receptive field that is manageable [62] and to maintain simplicity, we have opted to employ a single-layer GAT. Complementing this, we used a Multi-layer Perceptron (MLP) [63] with a three-layer configuration and RELU as the non-linear activation function to establish the neural network backbone of DDQN. The MLP’s output signifies the Q-value for the 2 distinct actions for GATDDQN. For further extension to scenarios like passenger transfer[5], relocations[15] and multi-hop [4], these actions could be further expanded to cover the decision of whether to refrain from picking up a potential new passenger or to indeed embark the passenger within one of the zones or stations on the 57-zone map. With the vehicle capacity set to three, the input state into the GATs constitutes a 1-by-14 tensor representing each vehicle’s state, while the final aggregated input state to DDQN is a 1-by-128 tensor. We set the memory capacity $C$ at 20,000 and the mini-batch size of $D$ as 1024. For training updates, further technical considerations included employing the Mean Squared Error (MSE) loss and utilizing Adam [64] as the optimizer for GATDDQN, with an assigned learning rate $\alpha=0.01$, soft update rate $\rho=0.005$, and gradient clipping threshold as 0.05. For graph sampling, we set the fixed number of agents as 30 and observation distance as equal to bipartite match distance $R_{match}$. Regarding exploration and exploitation trade-offs, we set the initial exploration rate $\epsilon$ as 1, exploration decay rate $\beta=0.996$, and exploration final value $\epsilon_{T}$ as 0.005. For every training in the following sessions, we standardized the comparison by configuring representative frameworks with neural network architectures and hyper-parameters that closely mirror those used in our BMG-Q model, and train our BMG-Q and representative frameworks for around 2000 episodes on Intel 14700K CPU and NVIDIA GEFORCE 4080 GPU desktop setup [^1] Although the training process may take up to two days to complete, our BMG-Q is efficient for real-time decision-making at scale. Once the BMG-Q is trained, for each minute of dispatch decisions involving approximately 1000 vehicles and 300 orders, the dispatch process takes only about 1 to 2 seconds on our desktop..
@@ VI-B Effectiveness and Scalability of BMG-Q Framework
Firstly, to test the effectiveness of BMG-Q framework in training, we compare our framework with three representative frameworks from modelling and MARL in ride-pooling: Greedy [3, 15], ILP + Independent RL [15, 4, 50, 16, 5, 6], ILP + Independent RL Considering Agents Nearby [38], and vanilla ILP + Attention-based MARL [25, 43]. For Greedy framework, we train a reward model offline till convergence and replace the Q function $Q(s,a)$ in Equation (11) with reward function $r(s,a)$. For ILP + Independent framework, we adopt ILPDDQN given in Algorithm 1 and also add exploration inside for the aim of fair comparisons. For ILP + Independent RL Considering Agents Nearby, we extend ILPDDQN baseline by incorporating the count of other agents and requests within its current zone into the agent’s state representations, termed as IQL_CAN. For vanilla ILP + Attention-based MARL, we remove our localized graph, gradient clipping and graph sampling strategy, which however lead the training to become significantly unstable in large-scale system with 1000 agents (thus not shown in the simulation results).
@@
Accumulative Total Reward (×$\times 10^{6}$) 1.007 0.925 0.901 0.900
@@ VI-C Robustness of BMG-Q Framework to Task Variations
In practice, the trained neural network may have to be applied to distinct scenarios, such as varying market conditions and fleet sizes. To validate the robustness of our BMG-Q framework under distinct scenarios, we first train the neural network on a specific scenario—peak hours on a Wednesday using a fleet of 1000 cars, and then test it across a range of fleet sizes and task variations. Specifically, we first explore its adaptability to different fleet size configurations of 800, 1000, and 1200 vehicles during the same time period, as presented in Table II. The results demonstrate that BMG-Q outperforms the ILPDDQN model consistently across multiple metrics, including rewards and order pickups, regardless of the fleet size. Subsequently, we extended our evaluation of the BMG-Q framework to check its performance across task variations over an entire month, as depicted in Figure 10. This comparison sheds light on the framework’s robustness against fluctuating operational conditions with 1000 vehicles. It was observed that the BMG-Q framework consistently outstripped both the ILPDDQN and Greedy baselines in terms of daily average rewards over the span of four weeks. From these two sets of validation exercises, we can infer that our BMG-Q framework demonstrates robustness compared to previous benchmarks not only to variations in the fleet size but also to task variations common in ride-pooling scenarios of TNCs, such as varying fleet sizes, and day-to-day policy adaptation [65, 66].
@@
Accumulative Total Reward (×$\times 10^{6}$) 1.000 1.010 0.980 0.974
@@ VI-D Sensitivity Analysis of BMG-Q
To evaluate the sensitivity of our proposed BMG-Q framework with respect to training hyperparameters, we test how the BMG-Q training performs under variations of critical hyperparameters for GATDDQN training. These parameters included the learning rate (lr), memory capacity, the number of samples in the graph sampling techniques, and the seat capacity of the ride-pooling vehicles.
First, we conducted a sensitivity analysis by retraining the model under four distinct learning rate settings: lr = 0.005, lr = 0.009, lr = 0.011, and lr = 0.02. We compared the training performances of these settings against our initial learning rate of lr = 0.01, prior to convergence. The results, depicted in Figure 11, demonstrate that the training performance of the BMG-Q framework remains consistently stable across these varied learning rate settings.
Figure 11: Sensitivity analaysis for BMG-Q under different learning rate in training
Second, we conducted another sensitivity analysis by retraining our model under two different experience memory capacities: Capacity = 10,000 and Capacity = 30,000. We compared the training performances of these settings against our initial experience memory setting of Capacity = 20,000, prior to convergence. The results, shown in Figure 12, indicate that the training performance of the BMG-Q framework remains consistently stable across the varied experience memory capacity settings.
Figure 12: Sensitivity analaysis for BMG-Q under different memory capacity in training
Third, we conducted a comparative analysis of the training performance when varying the number of neighboring vehicles sampled for each ego vehicle. Specifically, we trained the neural network on ride-hailing data from a Wednesday scenario involving 1000 cars, with the results illustrated in Figure 13. It is noteworthy that our training framework remains stable during the training phase in a large-scale system, thanks to our gradient clipping and graph sampling strategy, and this stability is maintained irrespective of the number of neighboring vehicles sampled. Furthermore, we interestingly discovered that once the number of sampled vehicles reaches a certain threshold, further increasing the sample size does not significantly impact the training performance (e.g., the brown curve vs. the purple curve), which validates that the use of localized bipartite graph while sampling a limited number of neighboring vehicles can well capture the interdependence between agents.
@@
VI-E Ablation Study of GAT in BMG-Q Framework
To further support our previous analysis in Section IV, we conduct an ablation study by modifying our graph aggregation method from GAT to two variants of GraphSAGE as described by Hamilton et al.[45]: GraphSAGE-Mean (Mean Aggregator) and GraphSAGE-Max (Max Aggregator). We retrained our system using each of these aggregation methods separately for the same task—managing a fleet of 1000 ride-pooling vehicles in Wednesday peak hour. The training plot is shown in Figure 14. From the figure, it is evident that our BMG-Q model, which utilizes GAT as the graph aggregator, exhibits notable improvements in performance relative to the other two baseline methods. Specifically, the GAT line (in purple) consistently achieves higher cumulative total rewards throughout the training process, compared to the lower and more variable trajectories observed with GraphSAGE_Mean (in orange) and GraphSAGE_Max (in blue). Furthermore, better interdependency encoding offered by GAT also leads to better training stability of our dispatch system, as indicated by the smaller confidence intervals in its reward trajectory.
Figure 14: Comparisons of BMG-Q with different graph aggregation strategies

### v2

@@ V-A Dynamic ILP via Posterior Score Function
In existing literature, the ILP for the bipartite matching process typically relies on $Q(s,a,\theta)$, such as (11) in Section II. However, this approach encounters two issues: (1) it lacks an exploration mechanism in the bipartite matching process, which restricts the exploration of the vehicle agents’ state space; (2) matching based solely on $Q(s,a,\theta)$ is prone to bias due to variability in the estimates, which can affect the decision-making process. These are particularly crucial in the case of ride-pooling vehicle dispatch, where ride-pooling requires agents now need to track detailed origin-destination (O-D) pairs and precise time detour information for each passenger in their states and thus have much larger state space compared with previous ride-sourcing setting. Therefore, sticking to conventional value function settings as in [9] will potentially lead to suboptimal solutions.
To deal with the above two issues and consider the importance of accurate value function [46], here we propose posterior score function to better integrate the value function of GATDDQN with ILP. The formulation of our posterior score function $S(s_{n,t},g_{n,t},a_{n,t})$ is given in Equation (26) below, with its visualization shown in Figure 3. To effectively balance exploration and exploitation, we implement an $\epsilon$-greedy strategy and introduce a term $S_{\text{explore}}$ during the exploration phase. This term represents the upper bound of the Q-value, $S(s_{n,t},g_{n,t},a_{n,t})$, which is set to a significantly high value (e.g., 100,000) to encourage exploration in exploration stage:
@@
Figure 3: Dynamic ILP with posterior score function. Our proposed posterior score function $S(s_{n,t},g_{n,t},a_{n,t})$ combines (1) an exploration-exploitation trade-off strategy with (2) GATDDQN-derived advantage estimation to enhance matching decisions throughout both training and validation phases.
@@ V-B Training GATDDQN for Large-Scale System
[…]
@@
The gradient clipping enforces the agents to update the policy within a region so as to not over-fit to recent experience tuples and guarantee to improve its policy. Through our further training and validation, we find that the two tricks not only help to stabilize the training process but also make the framework more robust to task variability and parameter change.
@@ V-C Summary of the proposed BMG-Q Framework
The training details of our BMG-Q framework could be found at Algorithm 2. Specifically, after initializing the simulator and GATDDQN in steps 1 through 4, we enter the training phase. To achieve a balance between exploitation and exploration, we perform exploration decay to exploration rate of the bipartite matching process in step 5. This gradual reduction in exploration rate is designed to transition the focus from exploration to exploitation as the learning advances. In steps 9 to 12, similar to the DDQN backbone of ILPDDQN, our GATDDDQN backbone adopts double networks, experience replay, and soft update. Thanks to the localized bipartite match graph topology, graph sampling, and gradient clipping introduced in steps 9 to 11, GATDDQN backbone manages to learn to capture the localized interdependence in very large-scale system with thousands of agents, thus leading to more effective assignment decisions of the overall BMG-Q framework.
@@
1: Simulator Initialization: OSRM Router Model[27], Number of Vehicles NN, Matching Distance $R_{match}$
@@
4:  Initialize: Episode Order Requirements, and Number of Vehicles NN
@@
11:    Sample NN experience tuples $(s,g,a,r,s^{\prime},g^{\prime})$ in $M$ as mini-batch $D$ and use Equation (28) and (29) to update $\theta$.
@@
Figure 4: Overview of the proposed BMG-Q framework for ride-pooling vehicles dispatch. For every decision round, initially users submit orders through a mobile application, and the system updates and sorts these orders alongside vehicle information. Taking into account long-term uncertainties including the intricate interdependence of agents, the GATDDQN network evaluates and dynamically assigns orders to suitable vehicles using ILP. After assignments, vehicle agents execute their actions, with experiences collected for subsequent learning phases of the GATDDQN network. Concurrently, the Route Guider updates routes and estimated times of arrival, which are communicated back to the central platform. (Part of icons and map are from [32, 11]).
@@ VI-A Simulation Setup
The simulation environment for this study is based on the public dataset of taxi trips in Manhattan, New York City [1, 38]. This dataset includes detailed information for each trip, such as pickup and dropoff times, origin and destination geo-coordinates, trip distance, and duration. Focusing on peak hours—specifically from 8:00 AM to 10:00 AM—we tailored the training dataset to include data from trips that occurred between 8:00 AM and 8:30 AM on May 4, 2016. During this half-hour period, the average order density reached approximately 275 trips per minute in central Manhattan, totaling about 8,250 orders. For the validation dataset, we similarly extracted data from trips within the same half-hour window but on various days throughout May 2016. We divided Manhattan into 57 zones. This zoning was informed by the distribution of orders and a resolution of 800m x 800m was used, for which a visualization is shown in Figure 5. To serve these demand with a minimum service rate of 85% across all RL methodologies tested in this study, we set the number of ride-pooling agents as 1000 and number of vacant seats of each vehicle as 3. To provide real-time route guidance and estimating the passengers’ onboard time, we employed the OSRM model [27] through docker as our router. The coefficients of reward function is set as ${\beta_{0}}=100$, ${\beta_{1}}=40$, ${\beta_{2}}=5$, ${\beta_{3}}=2$, ${\beta_{5}}=20$, thre=15{thre}=15, with the aim to encourage agents to pick up more orders but not result in large average detours of passengers. For bipartite match process, we set matching distance $R_{match}$ as 1.2 km and any requests that remain unmatched with vehicles for more than five minutes will be automatically rejected. Additionally, it’s important to note that our focus is on fully optimizing the potential of the ride-pooling fleet order dispatch during peak hours, when the demand is high and the occupancy of the vehicles are naturally high. Therefore, we have chosen not to include rebalancing operations [3] in our approaches or any of the benchmark approaches considered in this study. The incorporation of rebalancing operations is left for future work.
@@
In implementing GATDDQN, we adopted a linear transformation as message passing layer and multi-head attentions with a head count of 3 as aggregation layer to form the backbone structure of GATs. For the sake of a receptive field that is manageable [48] and to maintain simplicity, we have opted to employ a single-layer GAT. Complementing this, we used a Multi-layer Perceptron (MLP) with a three-layer configuration and RELU as the non-linear activation function to establish the neural network backbone of DDQN. The hidden unit of MLP is set as 256 and MLP’s output signifies the Q-value for the distinct actions of GATDDQN. For further extension to scenarios like passenger transfer[9] and relocations[1], these actions could be further expanded to cover the decision of whether to refrain from picking up a potential new passenger or to indeed embark the passenger within one of the zones or stations on the 57-zone map. With the vehicle capacity set to three, the input state into the GATs constitutes a 1-by-14 tensor representing each vehicle’s state, while the final aggregated input state to DDQN is a 1-by-128 tensor. We set the memory capacity $C$ at 20,000 and the mini-batch size of $D$ as 1024. For training updates, further technical considerations included employing the Mean Squared Error (MSE) loss and utilizing Adam as the optimizer for GATDDQN, with an assigned learning rate $\alpha=0.01$, soft update rate $\rho=0.005$, and gradient clipping threshold as 0.05. For graph sampling, we set the fixed number of agents as 30 and observation distance as equal to bipartite match distance $R_{match}$. Regarding exploration and exploitation trade-offs, we set the initial exploration rate $\epsilon$ as 1, exploration decay rate $\beta=0.996$, and exploration final value $\epsilon_{T}$ as 0.005. For every training in the following sessions, we standardized the comparison by configuring representative frameworks with neural network architectures and hyper-parameters that closely mirror those used in our BMG-Q model, and train our BMG-Q and representative frameworks for around 2000 episodes on Intel 14700K CPU and NVIDIA GEFORCE 4080 GPU desktop setup [^1] Although the training process may take up to two days to complete, our BMG-Q is efficient for real-time decision-making at scale. Once the BMG-Q is trained, for each minute of dispatch decisions involving approximately 1000 vehicles and 300 orders, the dispatch process takes only about 1 to 2 seconds on our desktop..
@@ VI-B Effectiveness and Scalability of BMG-Q Framework
Firstly, to test the effectiveness of BMG-Q framework in training, we compare our framework with three representative frameworks from modelling and MARL in ride-pooling: Greedy [3, 1], ILP + Independent RL [1, 31, 9, 43], ILP + Independent RL Considering Agents Nearby [33], and vanilla ILP + Attention-based MARL [20, 8]. For Greedy framework, we train a reward model offline till convergence and replace the Q function $Q(s,a)$ in Equation (11) with reward function $r(s,a)$. For ILP + Independent framework, we adopt ILPDDQN given in Algorithm 1 and also add exploration inside for the aim of fair comparisons. For ILP + Independent RL Considering Agents Nearby, we extend ILPDDQN baseline by incorporating the count of other agents and requests within its current zone into the agent’s state representations, termed as IQL_CAN. For vanilla ILP + Attention-based MARL, we remove our localized graph, gradient clipping and graph sampling strategy, which however lead the training to become significantly unstable in large-scale system with 1000 agents (thus not shown in the simulation results).
@@
Accumulative Total Reward ($\times 10^{6}$) 1.007 0.925 0.901 0.900
@@ VI-C Robustness of BMG-Q Framework to Task Variations
In practice, the trained neural network may have to be applied to distinct scenarios, such as varying market conditions and fleet sizes. To validate the robustness of our BMG-Q framework under distinct scenarios, we first train the neural network on a specific scenario—peak hours on a Wednesday using a fleet of 1000 cars, and then test it across a range of fleet sizes and task variations. Specifically, we first explore its adaptability to different fleet size configurations of 800, 1000, and 1200 vehicles during the same time period, as presented in Table II. The results demonstrate that BMG-Q outperforms the ILPDDQN model consistently across multiple metrics, including rewards and order pickups, regardless of the fleet size. Subsequently, we extended our evaluation of the BMG-Q framework to check its performance across task variations over an entire month, as depicted in Figure 9. This comparison sheds light on the framework’s robustness against fluctuating operational conditions with 1000 vehicles. It was observed that the BMG-Q framework consistently outstripped both the ILPDDQN and Greedy baselines in terms of daily average rewards over the span of four weeks. From these two sets of validation exercises, we can infer that our BMG-Q framework demonstrates robustness compared to previous benchmarks not only to variations in the fleet size but also to task variations common in ride-pooling scenarios of TNCs, such as varying fleet sizes, and day-to-day policy adaptation.
@@
Accumulative Total Reward ($\times 10^{6}$) 1.000 1.010 0.980 0.974
@@ VI-D Sensitivity Analysis of BMG-Q
To evaluate the sensitivity of our proposed BMG-Q framework with respect to training hyperparameters, we conducted a comparative analysis of the training performance when varying the number of neighboring vehicles sampled for each ego vehicle. Specifically, we trained the neural network on ride-hailing data from a Wednesday scenario involving 1000 cars, with the results illustrated in Figure 10. It is noteworthy that our training framework remains stable during the training phase in a large-scale system, thanks to our gradient clipping and graph sampling strategy, and this stability is maintained irrespective of the number of neighboring vehicles sampled. Furthermore, we interestingly discovered that once the number of sampled vehicles reaches a certain threshold, further increasing the sample size does not significantly impact the training performance (e.g., the brown curve vs. the purple curve), which validates that the use of localized bipartite graph while sampling a limited number of neighboring vehicles can well capture the interdependence between agents.